# Interval and $\ell$-interval rational parking functions

### v3

@@ 1. Introduction
Throughout, we let $\mathbb{N}\coloneqq\{0,1,2,3,\ldots\}$, +$\mathbb{N}^{+}\coloneqq\{1,2,3,\ldots\}$, and for $n\in\mathbb{N}^{+}$, we let $[n]\coloneqq\{1,2,\ldots,n\}$. We begin by recalling a definition of parking functions, as introduced by Konheim and Weiss [konheimOccupancyDisciplineApplications1966]. With $n\in\mathbb{N}^{+}$, a tuple $\alpha=(a_{1},a_{2},\ldots,a_{n})\in[n]^{n}$ is a parking function of length $n$ if its nondecreasing rearrangement $\alpha^{\prime}=(a_{1}^{\prime},a_{2}^{\prime},\ldots,a_{n}^{\prime})$ satisfies $a_{i}^{\prime}\leq i$ for all $i\in[n]$. For example, $\alpha=(3,1,1,3,4,5)$ is a parking function of length six as its nondecreasing rearrangement $\alpha^{\prime}=(1,1,3,3,4,5)$ satisfies the required inequality conditions. However, $\alpha=(2,3,3,3)$ is not a parking function, as its nondecreasing rearrangement $\alpha^{\prime}=(2,3,3,3)$ does not satisfy the inequality conditions: In particular, $a_{1}=2\nleq 1$. Let $\mathrm{PF}_{n}\subseteq[n]^{n}$ denote the set of parking functions of length $n$. Several authors have provided proofs that $|\mathrm{PF}_{n}|=(n+1)^{n-1}$, see [konheimOccupancyDisciplineApplications1966, pyke, Pollock].
@@
Parking functions have numerous applications and connections to many fields in mathematics, including computing volumes of flow polytopes, hyperplane arrangements, and the sorting algorithms [beck2014parking, Benedetti_2019, harris2023lucky]. Work on parking functions includes the study of discrete statistics [adeniran2023pattern, GesselSeo, Schumacher_Descents_PF], families of parking functions with particular desirable properties [Hanoi, flat_pf], and many generalizations of parking functions [colaric2020interval, knaples, assortments, MVP]. Our work is motivated by two generalizations of parking functions, which we refer to as “rational” parking functions and interval parking functions. We describe those next.
First we consider the situation in which there are $n\in\mathbb{N}^{+}$ cars and $m\geq n$ parking spots on a one-way street. In this case, a preference $\alpha$ is an element of $[m]^{n}$, and the set $\mathrm{PF}_{n,m}$ consists of the preferences allowing all $n$ cars to park under the standard parking procedure. This generalization of parking functions is known as $(m,n)$-parking functions. In the case of $n$ and $m$ being coprime, $(m,n)$-parking functions are referred to as rational parking functions, see [armstrong2016rational]. In the current work, we simply refer to $(m,n)$-parking functions as rational parking functions without requiring $m$ and $n$ to be coprime. These parking functions were enumerated during the onset of parking function studies [konheimOccupancyDisciplineApplications1966, Lemma 2] and, implicitly, in the statistical work of Pyke [pyke], giving
@@
Another generalization of parking functions are interval parking functions which describe the case where cars have an interval of the street in which they want to park. Formally, interval parking functions are defined by a pair $(\alpha,\beta)$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$, satisfy, for all $i\in[n]$, the inequality $a_{i}\leq b_{i}\leq n$. For $i\in[n]$, we call $a_{i}$ the initial preference of car $i$, $b_{i}$ the tolerance of car $i$ (as it is the highest numbered parking spot in which car $i$ would tolerate parking), and the set $\{a_{i},a_{i}+1,\ldots,b_{i}\}$ the preference interval for car $i$. Throughout, we refer to $\beta$ as the tolerance vector. If, for each $i\in[n]$, car $i$ is able to park in its preference interval following the standard parking procedure, then the pair $(\alpha,\beta)$ is called an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$. Colaric, DeMuse, Martin, and Yin established [colaric2020interval, Proposition 3.2]:
@@
Fubini rankings and unit interval parking functions have received much recent attention in the literature. In [unit_pf, Theorem 2.9], the authors give a complete characterization of unit interval parking functions. In [unit_perm], the authors establish that unit interval parking functions with $k$ unlucky cars (cars that do not park in their preference) are in bijection with the $k$-dimensional faces of the permutohedron of order $n$. In [boolean], the authors consider the set of unit Fubini rankings, which is the intersection of the set of unit interval parking functions of length $n$ and the set of Fubini rankings. Through the set of unit Fubini rankings and their parking interpretation, they establish a complete characterization and enumeration for the total number of Boolean intervals in the weak Bruhat order of the symmetric group on $n$ letters, and the total number of Boolean intervals of rank $k$ in the same. Motivated by these results, we establish the following.
@@
      The specialization of $m=n$ is provided in Theorem 5.2, recovering the fact that unit interval parking functions are enumerated by the Fubini numbers [Hadaway_unit_interval].
@@ 2.2. Interval parking functions
We begin by remarking that in [colaric2020interval], interval parking functions were only defined for $m=n$. In this setting, interval parking functions are characterized as a pair $(\alpha,\beta)\in\mathrm{PF}_{n}\times[n]^{n}$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$ satisfies $b_{i}\geq a_{i}$ for all $1\leq i\leq n$ [colaric2020interval, Proposition 3.2]. Given $\alpha\in\mathrm{PF}_{n}$ and $\beta\in[n]^{n}$, car $i$ enters the street and using the standard parking procedure, attempts to park between spots $a_{i}$ and $b_{i}$ (inclusive), which we refer to as its preferred interval. If all cars can park within their preferred interval utilizing the standard parking procedure, then the pair $(\alpha,\beta)$ is an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$ and recall (1.2) gives $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
@@ 2.3. Unit interval parking functions
Hadaway and Harris introduced unit interval parking functions, which are a subset of $\mathrm{PF}_{n}$ in which all cars park in their preferred spot or one spot further down the street [Hadaway_unit_interval]. Let $\mathrm{UPF}_{n}$ denote the set of unit interval parking functions of length $n$. Note, $(1,2,3,4,5)$, $(1,1,3,4,5)$, $(1,1,2,4,5)$ are elements of $\mathrm{UPF}_{5}$, whereas $(1,1,1,1,1)$ is not a unit interval parking function as car three cannot park in its preferred spot nor the one after it.
@@
[Hadaway_unit_interval, Theorem 5.12] If $n\geq 1$, then
@@
To establish Theorem 2.2, Hadaway and Harris proved that the set of unit interval parking functions is in bijection with the set of Fubini rankings ([Hadaway_unit_interval, Theorem 5.12]). As Fubini rankings play a role in our work, we define them next.
@@
A Fubini ranking of length $n$ is a tuple $r=(r_{1},r_{2},\ldots,r_{n})\in[n]^{n}$ that records a valid ranking of $n$ competitors with ties allowed (i.e., multiple competitors can be tied and have the same rank). If $k$ competitors are tied for rank $i$, the $k-1$ subsequent ranks $i+1,i+2,\ldots,i+k-1$ are disallowed. [^1] As noted in [boolean]: “One noteworthy instance of the tie condition of a Fubini ranking took place at the men’s high jump event at the Summer 2020 Olympics [olympics]. In this competition, Mutaz Essa Barshim of Qatar and Gianmarco Tamberi of Italy led the final round. Both athletes cleared 2.37 meters, but neither of them cleared 2.39 meters. Upon being presented the option of a “jump-off” to determine the sole winner, they agreed to instead share the gold medal. The next best rank was held by Maksim Nedasekau of Belarus, who obtained the bronze medal.”
For example, $(8,1,2,5,6,2,2,6)$ is Fubini ranking, in which competitor one ranks eighth, competitor two ranks first, competitors three, six, and seven tie at rank two, competitor four ranks fifth, and competitors five and eight rank sixth. Note $(1,1,1,2)$ is not a Fubini ranking as the triple tie for rank one would disallow ranks 2 and 3, making rank 4 the next available. We let $\text{FR}_{n}$ denote the set of Fubini rankings of length $n$. We remark that Cayley [cayley_2009] showed that $|\text{FR}_{n}|=\mathrm{Fub}_{n}$; this together with the bijection of Hadway and Harris, establishes (2.1).
@@ 3. Interval and $\ell$-interval rational parking functions
[colaric2020interval, Proposition 3.2] If $n\geq 1$, then $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
@@ 3.3. $\ell-$interval rational parking functions
We proceed as in the proof of Theorem 3.7 and assume that the last car parks in spot $z$. There are $m\geq n$ parking spots where $n$ is the number of cars. If car $n$ parks in spot $z$, there are $10\leq k\leq n-$ 1 cars parked to the left of spot $z$ and the remaining $n-1-k$ cars park to the right of spot $z$. There are $\binom{n-1}{k}$ ways to select the cars that park to the left of spot $z$. Before proceeding with the count, we need to consider that the preference of car $n$ which parks in spot $z$ is dependent on the number of contiguous cars which occupy spots immediately to the left of spot $z$, as then car $n$ could prefer some of those parking spots. Let $L$ be the number of cars parked immediately to the left of spot $z$ (from the $k$ cars parked somewhere to the left), with the previous spot empty. Note $0\leq L\leq k$ and those cars occupy spots $z-L,z-L+1,\ldots,z-1$, and spot $z-L-1$ is empty. From the $k$ cars which parked to the left of spot $z$, we must select $L$, which we can do in $\binom{k}{L}$ ways. The $L$ cars can have $|\mathrm{\mathrm{IPF}}_{L,L}(\ell)|=|\mathrm{\mathrm{IPF}}_{L}(\ell)|$ possible preferences so as to park in the spots required. Then the remaining $k-L$ cars must park in spots $21,2,\ldots,z-L-$ 2, as we must leave spot $z-L-1$ empty. Those cars can have $|\mathrm{\mathrm{IPF}}_{k-L,z-L-2}(\ell)|$ preferences. To ensure car $n$ parks in spot $z$, we note that the possible preferences depend on both $\ell$ and $L$. Namely, car $n$ can have $\min(\ell,L)+1$ preferences. Hence, the number of ways to park $k$ cars to the left of spot $z$ so as to satisfy all of these conditions, and so that car $n$ park in spot $z$, is given by
@@ 4.1. A connection to lattice paths
Following [armstrong2016rational], we recall that Dyck paths can be encoded via a Dyck word of length $2n$ consisting of $n$ letters $N$, representing north steps, and $n$ letters $E$, representing east steps, satisfying
@@
The enumeration of Dyck paths based on their exact height is well-known (OEIS A080936) and is credited to Grossman [Grossman], although we were unable to access the article to confirm this. Deutsch provides a generating function for the number of Dyck paths with semilength $2n$ and height $k$ (see OEIS A080936) as follows:
$h(k)=\frac{z^{k}}{f(k)\cdot f(k+1)},$
where $f(k)$ are polynomials defined by $f(k)=\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\binom{k-i}{i}\cdot(-z)^{i}$. From this, Deutsch states that the generating function for the number of Dyck paths of height at most $k$ is
$H(k)=\frac{f(k)}{f(k+1)}.$ (4.4)
@@
For $k\geq 0$, define $f(k)=\displaystyle\sum_{i=0}^{\lfloor\frac{k}{2}\rfloor}\binom{k-i}{i}\cdot(-z)^{i}$. Then the generating function for the number of nondecreasing $\ell$-interval parking functions of length $n$ is
$H(\ell+1)=\frac{f(\ell+1)}{f(\ell+2)}.$
We can extend this result naturally to $\ell$-interval rational parking functions, which are in bijection with rational $(n,m)$-Dyck paths (see [ceballos2018signature] for an overview of rational Dyck paths and other Catalan objects, and see [christensenGeneralizationParkingFunctions2020] for a relationship between $k$-Naples parking functions and $k$-lattice paths). Let us continue with some definitions and notation.
@@
Definition 4.12 (Pages 761-762 in [ceballos2018signature]).
@@ 4.2. The members of $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ that are also Fubini rankings
We now consider the intersection of nondecreasing $\ell$-interval parking functions which are also Fubini rankings (Definition 2.3). The motivation for considering these sets is the recent work of Elder, Harris, Kretschmann, and Martínez Mori, who enumerated unit Fubini rankings which are unit interval parking functions of length $n$ that are also Fubini rankings. They established the following enumerative result [boolean, Theorem 3.3].
@@
In Table 2, we give data for the cardinalities of the set $\mathrm{FR}_{n}^{\uparrow}(\ell)$ when $0\leq n\leq 11$ and $0\leq\ell\leq 6$. We direct interested readers to [milesGeneralizedFibonacciNumbers1960]).
@@
As noted at the start of this section, Elder, Harris, Kretschmann, and Martínez Mori studied the set $\mathrm{FR}_{n}(1)$ [boolean]. They proved that $|\mathrm{FR}_{n}(1)|$ was equal to the number of Boolean intervals in the weak order lattice of $\mathfrak{S}_{n}$ [boolean, Theorem 1.2]. It remains an open problem to determine if there is some correspondence between the set $\mathrm{FR}_{n}(\ell)$ and other structures in the weak order lattice of $\mathfrak{S}_{n}$.

### v4

@@ 1. Introduction
Throughout, we let $\mathbb{N}\coloneqq\{0,1,2,3,\ldots\}$, +$\mathbb{N}^{+}\coloneqq\{1,2,3,\ldots\}$, and for $n\in\mathbb{N}^{+}$, we let $[n]\coloneqq\{1,2,\ldots,n\}$. We begin by recalling a definition of parking functions, as introduced by Konheim and Weiss [24]. With $n\in\mathbb{N}^{+}$, a tuple $\alpha=(a_{1},a_{2},\ldots,a_{n})\in[n]^{n}$ is a parking function of length $n$ if its nondecreasing rearrangement $\alpha^{\prime}=(a_{1}^{\prime},a_{2}^{\prime},\ldots,a_{n}^{\prime})$ satisfies $a_{i}^{\prime}\leq i$ for all $i\in[n]$. For example, $\alpha=(3,1,1,3,4,5)$ is a parking function of length six as its nondecreasing rearrangement $\alpha^{\prime}=(1,1,3,3,4,5)$ satisfies the required inequality conditions. However, $\alpha=(2,3,3,3)$ is not a parking function, as its nondecreasing rearrangement $\alpha^{\prime}=(2,3,3,3)$ does not satisfy the inequality conditions: In particular, $a_{1}=2\nleq 1$. Let $\mathrm{PF}_{n}\subseteq[n]^{n}$ denote the set of parking functions of length $n$. Several authors have provided proofs that $|\mathrm{PF}_{n}|=(n+1)^{n-1}$, see [24, 28, 30].
@@
Parking functions have numerous applications and connections to many fields in mathematics, including computing volumes of flow polytopes, hyperplane arrangements, and the sorting algorithms [5, 6, 23]. Work on parking functions includes the study of discrete statistics [1, 18, 29], families of parking functions with particular desirable properties [2, 16], and many generalizations of parking functions [12, 13, 17, 22]. Our work is motivated by two generalizations of parking functions, which we refer to as “rational” parking functions and interval parking functions. We describe those next.
First we consider the situation in which there are $n\in\mathbb{N}^{+}$ cars and $m\geq n$ parking spots on a one-way street. In this case, a preference $\alpha$ is an element of $[m]^{n}$, and the set $\mathrm{PF}_{n,m}$ consists of the preferences allowing all $n$ cars to park under the standard parking procedure. This generalization of parking functions is known as $(m,n)$-parking functions. In the case of $n$ and $m$ being coprime, $(m,n)$-parking functions are referred to as rational parking functions, see [4]. In the current work, we simply refer to $(m,n)$-parking functions as rational parking functions without requiring $m$ and $n$ to be coprime. These parking functions were enumerated during the onset of parking function studies [24, Lemma 2] and, implicitly, in the statistical work of Pyke [28], giving
@@
Another generalization of parking functions are interval parking functions which describe the case where cars have an interval of the street in which they want to park. Formally, interval parking functions are defined by a pair $(\alpha,\beta)$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$, satisfy, for all $i\in[n]$, the inequality $a_{i}\leq b_{i}\leq n$. For $i\in[n]$, we call $a_{i}$ the initial preference of car $i$, $b_{i}$ the tolerance of car $i$ (as it is the highest numbered parking spot in which car $i$ would tolerate parking), and the set $\{a_{i},a_{i}+1,\ldots,b_{i}\}$ the preference interval for car $i$. Throughout, we refer to $\beta$ as the tolerance vector. If, for each $i\in[n]$, car $i$ is able to park in its preference interval following the standard parking procedure, then the pair $(\alpha,\beta)$ is called an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$. Colaric, DeMuse, Martin, and Yin established [12, Proposition 3.2]:
@@
Fubini rankings and unit interval parking functions have received much recent attention in the literature. In [7, Theorem 2.9], the authors give a complete characterization of unit interval parking functions. In [10], the authors establish that unit interval parking functions with $k$ unlucky cars (cars that do not park in their preference) are in bijection with the $k$-dimensional faces of the permutohedron of order $n$. In [15], the authors consider the set of unit Fubini rankings, which is the intersection of the set of unit interval parking functions of length $n$ and the set of Fubini rankings. Through the set of unit Fubini rankings and their parking interpretation, they establish a complete characterization and enumeration for the total number of Boolean intervals in the weak Bruhat order of the symmetric group on $n$ letters, and the total number of Boolean intervals of rank $k$ in the same. Motivated by these results, we establish the following.
@@
      The specialization of $m=n$ is provided in Theorem 5.2, recovering the fact that unit interval parking functions are enumerated by the Fubini numbers [21].
@@ 2.2. Interval parking functions
We begin by remarking that in [12], interval parking functions were only defined for $m=n$. In this setting, interval parking functions are characterized as a pair $(\alpha,\beta)\in\mathrm{PF}_{n}\times[n]^{n}$, where $\alpha=(a_{1},a_{2},\ldots,a_{n})\in\mathrm{PF}_{n}$ and $\beta=(b_{1},b_{2},\ldots,b_{n})\in[n]^{n}$ satisfies $b_{i}\geq a_{i}$ for all $1\leq i\leq n$ [12, Proposition 3.2]. Given $\alpha\in\mathrm{PF}_{n}$ and $\beta\in[n]^{n}$, car $i$ enters the street and using the standard parking procedure, attempts to park between spots $a_{i}$ and $b_{i}$ (inclusive), which we refer to as its preferred interval. If all cars can park within their preferred interval utilizing the standard parking procedure, then the pair $(\alpha,\beta)$ is an interval parking function. We let $\mathrm{\mathrm{IPF}}_{n}$ denote the set of interval parking functions of length $n$ and recall (1.2) gives $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
@@ 2.3. Unit interval parking functions
Hadaway and Harris introduced unit interval parking functions, which are a subset of $\mathrm{PF}_{n}$ in which all cars park in their preferred spot or one spot further down the street [21]. Let $\mathrm{UPF}_{n}$ denote the set of unit interval parking functions of length $n$. Note, $(1,2,3,4,5)$, $(1,1,3,4,5)$, $(1,1,2,4,5)$ are elements of $\mathrm{UPF}_{5}$, whereas $(1,1,1,1,1)$ is not a unit interval parking function as car three cannot park in its preferred spot nor the one after it.
@@
[21, Theorem 5.12] If $n\geq 1$, then
@@
To establish Theorem 2.2, Hadaway and Harris proved that the set of unit interval parking functions is in bijection with the set of Fubini rankings ([21, Theorem 5.12]). As Fubini rankings play a role in our work, we define them next.
@@
A Fubini ranking of length $n$ is a tuple $r=(r_{1},r_{2},\ldots,r_{n})\in[n]^{n}$ that records a valid ranking of $n$ competitors with ties allowed (i.e., multiple competitors can be tied and have the same rank). If $k$ competitors are tied for rank $i$, the $k-1$ subsequent ranks $i+1,i+2,\ldots,i+k-1$ are disallowed. [^1] As noted in [15]: “One noteworthy instance of the tie condition of a Fubini ranking took place at the men’s high jump event at the Summer 2020 Olympics [14]. In this competition, Mutaz Essa Barshim of Qatar and Gianmarco Tamberi of Italy led the final round. Both athletes cleared 2.37 meters, but neither of them cleared 2.39 meters. Upon being presented the option of a “jump-off” to determine the sole winner, they agreed to instead share the gold medal. The next best rank was held by Maksim Nedasekau of Belarus, who obtained the bronze medal.”
For example, $(8,1,2,5,6,2,2,6)$ is Fubini ranking, in which competitor one ranks eighth, competitor two ranks first, competitors three, six, and seven tie at rank two, competitor four ranks fifth, and competitors five and eight rank sixth. Note $(1,1,1,2)$ is not a Fubini ranking as the triple tie for rank one would disallow ranks 2 and 3, making rank 4 the next available. We let $\text{FR}_{n}$ denote the set of Fubini rankings of length $n$. We remark that Cayley [8] showed that $|\text{FR}_{n}|=\mathrm{Fub}_{n}$; this together with the bijection of Hadway and Harris, establishes (2.1).
@@ 3. Interval and $\ell$-interval rational parking functions
[12, Proposition 3.2] If $n\geq 1$, then $|\mathrm{\mathrm{IPF}}_{n}|=n!(n+1)^{n-1}$.
@@ 3.3. $\ell-$interval rational parking functions
We proceed as in the proof of Theorem 3.7 and assume that the last car parks in spot $z$. There are $m\geq n$ parking spots where $n$ is the number of cars. If car $n$ parks in spot $z$, there are $10\leq k\leq z-$ 1 cars parked to the left of spot $z$ and the remaining $n-1-k$ cars park to the right of spot $z$. There are $\binom{n-1}{k}$ ways to select the cars that park to the left of spot $z$. Before proceeding with the count, we need to consider that the preference of car $n$ which parks in spot $z$ is dependent on the number of contiguous cars which occupy spots immediately to the left of spot $z$, as then car $n$ could prefer some of those parking spots. Let $L$ be the number of cars parked immediately to the left of spot $z$ (from the $k$ cars parked somewhere to the left), with the previous spot empty. Note $0\leq L\leq k$ and those cars occupy spots $z-L,z-L+1,\ldots,z-1$, and spot $z-L-1$ is empty. From the $k$ cars which parked to the left of spot $z$, we must select $L$, which we can do in $\binom{k}{L}$ ways. The $L$ cars can have $|\mathrm{\mathrm{IPF}}_{L,L}(\ell)|=|\mathrm{\mathrm{IPF}}_{L}(\ell)|$ possible preferences so as to park in the spots required. Then the remaining $k-L$ cars must park in spots $21,2,\ldots,z-L-$ 2, as we must leave spot $z-L-1$ empty. Those cars can have $|\mathrm{\mathrm{IPF}}_{k-L,z-L-2}(\ell)|$ preferences. To ensure car $n$ parks in spot $z$, we note that the possible preferences depend on both $\ell$ and $L$. Namely, car $n$ can have $\min(\ell,L)+1$ preferences. Hence, the number of ways to park $k$ cars to the left of spot $z$ so as to satisfy all of these conditions, and so that car $n$ park in spot $z$, is given by
@@ 4.1. A connection to lattice paths
Following [4], we recall that Dyck paths can be encoded via a Dyck word of length $2n$ consisting of $n$ letters $N$, representing north steps, and $n$ letters $E$, representing east steps, satisfying
@@
The enumeration of Dyck paths based on their exact height is well-known (OEIS A080936) and is credited to Grossman [20], although we were unable to access the article to confirm this. Thanks to a comment from Emeric Deutsch posted on OEIS A080936 we found the following: By [25, Page 37-38] the number of Dyck paths with semilength $2n$ and height at most $k$, denoted $H(k)$, is given by
$H(k)=[t^{k}]\left(\frac{f_{k}(t)}{f_{k+1}(t)}\right),$ (4.4)
which is the coefficient of $t^{k}$ in the generating function $f_{k}(t)/f_{k+1}(t)$, where $f_{0}(t)=f_{1}(t)=1$ and $f_{k+1}(t)=f_{k}(t)-t\cdot f_{k-1}(t)$. Note that the recurrence relation for the polynomials $f_{k}(t)$ was given in [25, Page 36, Equation 14].
@@
For $k\geq 0$, the number of nondecreasing $\ell$-interval parking functions of length $n$ is $H(\ell+1)$, as defined in Equation (4.4).
We can extend this result naturally to $\ell$-interval rational parking functions, which are in bijection with rational $(n,m)$-Dyck paths (see [9] for an overview of rational Dyck paths and other Catalan objects, and see [11] for a relationship between $k$-Naples parking functions and $k$-lattice paths). Let us continue with some definitions and notation.
@@
Definition 4.12 (Pages 761-762 in [9]).
@@ 4.2. The members of $\mathrm{\mathrm{IPF}}_{n}^{\uparrow}(\ell)$ that are also Fubini rankings
We now consider the intersection of nondecreasing $\ell$-interval parking functions which are also Fubini rankings (Definition 2.3). The motivation for considering these sets is the recent work of Elder, Harris, Kretschmann, and Martínez Mori, who enumerated unit Fubini rankings which are unit interval parking functions of length $n$ that are also Fubini rankings. They established the following enumerative result [15, Theorem 3.3].
@@
In Table 2, we give data for the cardinalities of the set $\mathrm{FR}_{n}^{\uparrow}(\ell)$ when $0\leq n\leq 11$ and $0\leq\ell\leq 6$. We direct interested readers to [26]).
@@
As noted at the start of this section, Elder, Harris, Kretschmann, and Martínez Mori studied the set $\mathrm{FR}_{n}(1)$ [15]. They proved that $|\mathrm{FR}_{n}(1)|$ was equal to the number of Boolean intervals in the weak order lattice of $\mathfrak{S}_{n}$ [15, Theorem 1.2]. It remains an open problem to determine if there is some correspondence between the set $\mathrm{FR}_{n}(\ell)$ and other structures in the weak order lattice of $\mathfrak{S}_{n}$.